# An Introduction to Decentralized Training and Execution in Cooperative Multi-Agent Reinforcement Learning

### v3

@@
An Introduction to
Decentralized Training and Execution in
@@
Multi-agent reinforcement learning (MARL) has exploded in popularity in recent years. Many approaches have been developed but they can be divided into three main types: centralized training and execution (CTE), centralized training for decentralized execution (CTDE), and decentralized training and execution (DTE).
@@
CTE methods assume centralization during training and execution (e.g., with fast, free and perfect communication) and have the most information during execution. That is, the actions of each agent can depend on the information from all agents. As a result, a simple form of CTE can be achieved by using a single-agent RL method with the centralized action and observation spaces (maintaining a centralized action-observation history for the partially observable case). CTE methods can potentially outperform the decentralized execution methods (since they allow centralized control) but are less scalable as the (centralized) action and observation spaces scale exponentially with the number of agents. CTE is typically only used in the cooperative MARL case since centralized control implies cooperation. CTDE methods are perhaps the most common as they can use centralized information during training but execute in a decentralized manner—using only information available to that agent during execution. As a result, they can be more scalable than CTE methods, do not require communication during execution, and can often perform well. CTDE fits most naturally with the cooperative case, but can be potentially applied in competitive or mixed settings depending on what information is assumed to be observed. Decentralized training and execution methods make the fewest assumptions and are often simple to implement. In fact, as I’ll discuss, any single-agent RL method can be used for DTE by just letting each agent learn separately. Of course, there are pros and cons to such approaches as we discuss below. It is worth noting that DTE is required if no offline coordination is available. That is, if all agents must learn during online interactions without prior coordination, learning and execution must both be decentralized. DTE methods can be applied in cooperative, competitive, or mixed cases but this text will focus on the cooperative MARL case.
MARL methods can be further broken up into value-based and policy gradient methods. Value-based methods (e.g., Q-learning) learn a value function and then choose actions based on those values. Policy gradient methods learn an explicit policy representation and attempt to improve the policy in the direction of the gradient. Both classes of methods are widely used in MARL.
@@
In this text, I will first give a brief description of the cooperative MARL problem in the form of the Dec-POMDP. Then, I will discuss value-based DTE methods starting with independent Q-learning and its extensions and then discuss the extension to the deep case with DQN, the additional complications this causes, and methods that have been developed to (attempt to) address these issues. Next, I will discuss policy gradient DTE methods starting with independent REINFORCE (i.e., vanilla policy gradient), and then extending to the actor-critic case and deep variants (such as independent PPO). Finally, I will discuss some general topics related to DTE and future directions.
@@
The basics of reinforcement learning (in the single-agent setting) are not presented in this text. Anyone interested in RL should read the book by Sutton and Barto (2018). Similarly, for a broader overview of MARL, the recent book by Albrecht, Christianos and Schäfer is recommended (Albrecht et al., 2024).
It is worth noting that this is an introduction to the field of decentralized, cooperative MARL. This subarea has less work than subareas such as CTDE but it is still more extensive than presented here. I have included work that I believe is important for understanding the main concepts in the subarea and apologize for those that I have omitted.
1 The cooperative MARL problem: The Dec-POMDP
Figure 1: A depiction of cooperative MARL—a Dec-POMDP.
@@ 1 The cooperative MARL problem: The Dec-POMDP
The cooperative multi-agent reinforcement learning (MARL) problem can be represented as a Dec-POMDP (Oliehoek and Amato, 2016; Bernstein et al., 2002). Dec-POMDPs generalize POMDPs (Kaelbling et al., 1998) (and MDPs (Puterman, 1994)) to the multi-agent, decentralized setting. As depicted in Figure 1, multiple agents operate under uncertainty based on partial views of the world, with execution unfolding over time. At each step, every agent chooses an action (in parallel) based purely on locally observable information, resulting in each agent obtaining an observation and the team obtaining a joint reward. The shared reward function makes the problem cooperative, but their local views mean that execution is decentralized.
@@
A solution to a Dec-POMDP is a joint policy, denoted $\mathbf{\pi}$—a set of policies, one for each agent, each of which is denoted ${\pi}_{i}$. Because the state is not directly observed, it is typically beneficial for each agent to remember a history of its observations. Then, a local policy, ${\pi}_{i}$, for an agent is a mapping from local action-observation histories to actions, ${\mathbb{H}}_{i}\to{\mathbb{A}}_{i}$, where ${\mathbb{H}}_{i}$ is the set of local observation histories, ${h}_{i}=\{{a}_{i,1},{o}_{i,1},\ldots,{a}_{i,t},{o}_{i,t}\}$ [^3] Sometimes, ${o}_{i,0}$ is also included as an observation generated by the initial state distribution., by agent $i$ up to the current time step, $t$. Many researchers just use observation histories (without including actions), which is sufficient for deterministic policies but may not be for stochastic policies (Oliehoek and Amato, 2016). Note that histories have implicit time steps due to their length which we do not include in the notation (i.e., we always assume a history starts on the first time step and the last time step is defined by the number of action-observation pairs). We can denote the histories for all agents at a given time step as $\mathbf{h}{}=\langle{h}_{1},\ldots,{h}_{n}\rangle$. Because one policy is generated for each agent and these policies depend only on local observations, they operate in a decentralized manner.
While there always exists an optimal deterministic joint policy in Dec-POMDPs, we can also use stochastic joint policies, overloading notation with: $\mathbf{\pi}(\mathbf{a}|\mathbf{h}{})=\prod_{i\in\mathbb{I}}{\pi}_{i}({a}_{i}|{h}_{i})$, where ${\pi}_{i}({a}_{i}|{h}_{i})$ represents the local policy for agent $i$ and the probability of choosing action ${a}_{i}$ in history ${h}_{i}$. Deterministic policies will be used in the value-based methods in Section 2 while stochastic policies will be used in the policy gradient gradient methods in Section 3.
@@
$V^{\mathbf{\pi}}(\mathbf{h}{})=\sum_{s}P(s|\mathbf{h}{},{{{b}}_{0}})\Big{[}R(s,\mathbf{\pi}(\mathbf{h}{}))+\gamma\sum_{s^{\prime}}P(s^{\prime}|\mathbf{\pi}(\mathbf{h}{}),s)\sum_{\mathbf{o}}P(\mathbf{o}|\mathbf{\pi}(\mathbf{h}{}),s^{\prime})V^{\mathbf{\pi}}(\mathbf{h}{}\mathbf{\pi}(\mathbf{h}{})\mathbf{o})\Big{]}$
where $P(s|\mathbf{h}{},{{{b}}_{0}})$ is the probability of state $s$ after observing joint history $\mathbf{h}{}$ starting from state distribution ${{{b}}_{0}}$ and $\mathbf{\pi}(\mathbf{h}{})$ is the joint action taken at the joint history. Also, $\mathbf{h}{}\mathbf{\pi}(\mathbf{h}{})\mathbf{o}$ represents $\mathbf{h}{}^{\prime}$, the joint history after taking joint action $\mathbf{\pi}(\mathbf{h}{})$ in joint history $\mathbf{h}{}$ and observing joint observation $\mathbf{o}$. In the RL context, algorithms do not iterate over states and observations to explicitly calculate this expectation but approximate it through sampling. For the finite-horizon case, $V^{\mathbf{\pi}}(\mathbf{h}{})=0$ when the length of $\mathbf{h}{}$ equals the horizon $\mathcal{H}$, showing the value function includes the time step from the history.
We will also evaluate policies starting from the initial, null history as $V^{\mathbf{\pi}}(\mathbf{h}{}_{0})$. An optimal joint policy beginning at $\mathbf{h}{}_{0}$ is $\mathbf{\pi}^{*}(\mathbf{h}{}_{0})=\operatorname*{arg\,\!max}_{\mathbf{\pi}}V^{\mathbf{\pi}}(\mathbf{h}{}_{0})$. That is, the optimal joint policy is the set of local policies for each agent that provides the highest value, which is denoted $V^{*}$.
Reinforcement learning methods often use history-action values, $Q(\mathbf{h}{},\mathbf{a})$, rather than just history values $V(\mathbf{h}{})$. $Q^{\mathbf{\pi}}(\mathbf{h}{},\mathbf{a})$ is the value of choosing joint action $\mathbf{a}$ at joint history $\mathbf{h}{}$ and then continuing with policy $\mathbf{\pi}$,
$Q^{\mathbf{\pi}}(\mathbf{h}{},\mathbf{a})=\sum_{s}P(s|\mathbf{h}{},{{{b}}_{0}})\Big{[}R(\mathbf{a},s)+\gamma\sum_{s^{\prime}}P(s^{\prime}|\mathbf{a},s)\sum_{\mathbf{o}}P(\mathbf{o}|\mathbf{a},s^{\prime})Q^{\mathbf{\pi}}\big{(}\mathbf{h}{}\mathbf{a}\mathbf{o},\mathbf{\pi}(\mathbf{h}{}\mathbf{a}\mathbf{o})\big{)}\Big{]},$
while $Q^{*}(\mathbf{h}{},\mathbf{a})$ is the value of choosing action $\mathbf{a}$ at history $\mathbf{h}{}$ and then continuing with the optimal policy, $\mathbf{\pi}^{*}$.

### v4

# A First Introduction to Cooperative Multi-Agent Reinforcement Learning

@@
A First Introduction to
@@ 1.1 Overview
Multi-agent reinforcement learning (MARL) has exploded in popularity in recent years. While numerous approaches have been developed, they can be broadly categorized into three main types: centralized training and execution (CTE), centralized training for decentralized execution (CTDE), and decentralized training and execution (DTE).
@@
CTE methods assume centralization during training and execution (e.g., with fast, free, and perfect communication) and have the most information during execution. That is, the actions of each agent can depend on the information from all agents. As a result, a simple form of CTE can be achieved by using a single-agent RL method with centralized action and observation spaces (maintaining a centralized action-observation history for the partially observable case). CTE methods can potentially outperform the decentralized execution methods (since they allow centralized control) but are less scalable as the (centralized) action and observation spaces scale exponentially with the number of agents. CTE is typically only used in the cooperative MARL case since centralized control implies coordination on what actions will be selected by each agent.
CTDE methods are the most common, as they leverage centralized information during training while enabling decentralized execution—using only information available to that agent during execution. CTDE is the only paradigm that requires a separate training phase where any available information (e.g., other agent policies, underlying states) can be used. As a result, they can be more scalable than CTE methods, do not require communication during execution, and can often perform well. CTDE fits most naturally with the cooperative case, but may also apply in competitive or mixed settings depending on what information is assumed to be observed.
Decentralized training and execution methods make the fewest assumptions and are often simple to implement. In fact, any single-agent RL method can be used for DTE by just letting each agent learn separately. Naturally, these approaches have their respective advantages and disadvantages, which will be discussed. It is worth noting that DTE is required if no centralized training phase is available (e.g., through a centralized simulator), requiring all agents to learn during online interactions without prior coordination. DTE methods can be applied in cooperative, competitive, or mixed cases. MARL methods can be further categorized into value-based and policy gradient methods. Value-based methods (e.g., Q-learning) learn a value function and then choose actions based on those values. Policy gradient methods learn an explicit policy representation and attempt to improve the policy in the direction of the gradient. Both classes of methods are widely used in MARL.
This text is an introduction to cooperative MARL—MARL in which all agents share a single, joint reward. It is meant to explain the setting, basic concepts, and common methods for the CTE, CTDE, and DTE settings. It does not cover all work in cooperative MARL as the area is quite extensive. I have included work that I believe is important for understanding the main concepts in the area and apologize to those that I have omitted.
I will first give a brief description of the cooperative MARL problem in the form of the Dec-POMDP as well as a short background on the (single-agent) reinforcement learning methods that are relevant to this text. Then, I will discuss the centralized, decentralized, and CTDE settings and algorithms. Note that all the presented methods are model-free (i.e., they do not learn a model of the dynamics or observations) since the vast majority of cooperative MARL methods are model-free.
For the centralized case, I will discuss the setting as well as ways of modeling the CTE problem, simple applications of single-agent methods as well as some more scalable methods that exploit the multi-agent structure.
@@
For the decentralized case, I will give an overview of the DTE setting and then I will discuss the relevant value-based and policy gradient methods. I’ll start with value-based DTE methods such as independent Q-learning and its extensions and then I will also discuss the extension to the deep case with DQN, the additional complications this causes, and methods that have been developed to (attempt to) address these issues. Next, I will discuss policy gradient DTE methods starting with independent REINFORCE (Peshkin et al., 2000) (i.e., vanilla policy gradient), and then extending to the actor-critic case and deep variants (such as independent PPO (de Witt et al., 2020; Yu et al., 2022)). Finally, I will discuss some general topics related to DTE and future directions.
Finally, I will present an overview of CTDE and the two main classes of CTDE methods: value function factorization methods and centralized critic actor-critic methods. Value function factorization methods include the well-known VDN (Sunehag et al., 2017), QMIX (Rashid et al., 2018), and QPLEX (Wang et al., 2021a) approaches, while centralized critic methods include MADDPG (Lowe et al., 2017), COMA (Foerster et al., 2018b), and MAPPO (Yu et al., 2022). Lastly, I discuss other forms of CTDE such as adding centralized information to decentralized (i.e., independent) learners (such as parameter sharing) and decentralizing centralized solutions.
@@
A detailed treatment of reinforcement learning (in the single-agent setting) is not presented in this text. Anyone interested in RL should read the book by Sutton and Barto (2018). Similarly, for a broader overview of MARL, the recent book by Albrecht, Christianos and Schäfer is recommended (Albrecht et al., 2024).
1.2 The cooperative MARL problem: The Dec-POMDP
Figure 1.1: A depiction of cooperative MARL—a Dec-POMDP.
@@ 1.2 The cooperative MARL problem: The Dec-POMDP
The cooperative multi-agent reinforcement learning (MARL) problem can be represented as a decentralized partially observable Markov decision process (Dec-POMDP) (Oliehoek and Amato, 2016; Bernstein et al., 2002). Dec-POMDPs generalize POMDPs (Kaelbling et al., 1998) (and MDPs (Puterman, 1994)) to the multi-agent, decentralized setting. They are also a subclass of partially observable stochastic games (POSGs) where all the agents have the same reward (i.e., a cooperative POSG). As depicted in Figure 1.1, multiple agents operate under uncertainty based on partial views of the world, with execution unfolding over time. At each step, every agent chooses an action (in parallel) based purely on locally observable information, resulting in each agent obtaining an observation and the team obtaining a joint reward. The shared reward function makes the problem cooperative, but their local views mean that execution is decentralized.
@@
A solution to a Dec-POMDP is a joint policy, denoted $\mathbf{\pi}$—a set of policies, one for each agent, each of which is denoted ${\pi}_{i}$. Because the state is not directly observed, it is typically beneficial for each agent to remember a history of its observations (and potentially actions). Then, a local policy, ${\pi}_{i}$, for an agent is a mapping from local action-observation histories to actions or probability distributions over actions. A local deterministic policy for agent $i$, ${\pi}_{i}({h}_{i})={a}_{i}$, maps ${\mathbb{H}}_{i}\to{\mathbb{A}}_{i}$, where ${\mathbb{H}}_{i}$ is the set of local observation histories, ${h}_{i}=\{{a}_{i,0},{o}_{i,0},\ldots,{a}_{i,t-1},{o}_{i,t-1}\}$ [^3] Sometimes, ${o}_{i,0}$ is also included as an observation generated by the initial state distribution, by agent $i$ up to the current time step, $t$. Note that histories have implicit time steps due to their length which I do not include in the notation (i.e., I always assume a history starts on the first time step and the last time step is defined by the number of action-observation pairs). We can denote the joint histories for all agents at a given time step as $\mathbf{h}{}=\langle{h}_{1},\ldots,{h}_{n}\rangle$. A joint deterministic policy is denoted $\pi(\mathbf{h}{})=\mathbf{a}=\langle{\pi}_{1}({h}_{1}),\dots,{\pi}_{n}({h}_{n})\rangle=\langle{a}_{1},\ldots,{a}_{n}\rangle$. A stochastic local policy for agent $i$ is ${\pi}_{i}({a}_{i}|{h}_{i})$, representing the probability of choosing action ${a}_{i}$ in history ${h}_{i}$. A joint stochastic policy is denoted $\mathbf{\pi}(\mathbf{a}|\mathbf{h}{})=\prod_{i\in\mathbb{I}}{\pi}_{i}({a}_{i}|{h}_{i})$. Because one policy is generated for each agent and these policies depend only on local observations, they operate in a decentralized manner.
Many researchers just use observation histories (without including actions), which is sufficient for deterministic policies but may not be for stochastic policies (Oliehoek and Amato, 2016). Deterministic policies will be used in the value-based methods while stochastic policies will be used in the policy gradient gradient methods. There always exists an optimal deterministic joint policy in Dec-POMDPs (Oliehoek and Amato, 2016), but stochastic (or continuous) policies are needed for policy gradient methods.
@@
$\mathbf{V}^{\mathbf{\pi}}(\mathbf{h}{})=\sum_{s}P(s|\mathbf{h}{},{{{b}}_{0}})\Big{[}R(s,\mathbf{a})+\gamma\sum_{s^{\prime}}P(s^{\prime}|\mathbf{a},s)\sum_{\mathbf{o}}P(\mathbf{o}|\mathbf{a},s^{\prime})\mathbf{V}^{\mathbf{\pi}}(\mathbf{h}{}\mathbf{a}\mathbf{o})\Big{]}\Big{|}_{\mathbf{a}=\mathbf{\pi}(\mathbf{h}{})}$ (1.1)
where $P(s|\mathbf{h}{},{{{b}}_{0}})$ is the probability of state $s$ after observing joint history $\mathbf{h}{}$ starting from state distribution ${{{b}}_{0}}$ and $\mathbf{a}=\mathbf{\pi}(\mathbf{h}{})$ is the joint action taken at the joint history. Also, $\mathbf{h}{}\mathbf{a}\mathbf{o}$ represents $\mathbf{h}{}^{\prime}$, the joint history after taking joint action $\mathbf{a}$ in joint history $\mathbf{h}{}$ and observing joint observation $\mathbf{o}$. In the RL context, algorithms do not iterate over states and observations to explicitly calculate this expectation but approximate it through sampling. For the finite-horizon case, $\mathbf{V}^{\mathbf{\pi}}(\mathbf{h}{})=0$ when the length of $\mathbf{h}{}$ equals the horizon $\mathcal{H}$, showing the value function includes the time step from the history.
We will often want to evaluate policies starting from the beginning—starting at the initial state distribution before any action or observation. Starting from this initial, null history I denote the value of a policy as $\mathbf{V}^{\mathbf{\pi}}(\mathbf{h}{}_{0})$. An optimal joint policy beginning at $\mathbf{h}{}_{0}$ is $\operatorname*{arg\,\!max}_{\mathbf{\pi}}\mathbf{V}^{\mathbf{\pi}}(\mathbf{h}{}_{0})$, where the $\operatorname*{arg\,\!max}$ here denotes enumeration over decentralized policies. The optimal joint policy is then the set of local policies for each agent that provides the highest value, which is denoted $\mathbf{V}^{*}$.
Reinforcement learning methods often use history-action values, $\mathbf{Q}(\mathbf{h}{},\mathbf{a})$, rather than just history values $\mathbf{V}(\mathbf{h}{})$. $\mathbf{Q}^{\mathbf{\pi}}(\mathbf{h}{},\mathbf{a})$ is the value of choosing joint action $\mathbf{a}$ at joint history $\mathbf{h}{}$ and then continuing with policy $\mathbf{\pi}$,
$\mathbf{Q}^{\mathbf{\pi}}(\mathbf{h}{},\mathbf{a})=\sum_{s}P(s|\mathbf{h}{},{{{b}}_{0}})\Big{[}R(\mathbf{a},s)+\gamma\sum_{s^{\prime}}P(s^{\prime}|\mathbf{a},s)\sum_{\mathbf{o}}P(\mathbf{o}|\mathbf{a},s^{\prime})\mathbf{Q}^{\mathbf{\pi}}\big{(}\mathbf{h}{}\mathbf{a}\mathbf{o},\mathbf{a}^{\prime})|_{\mathbf{a}^{\prime}=\mathbf{\pi}(\mathbf{h}{}\mathbf{a}\mathbf{o})}\big{)}\Big{]},$ (1.2)
while $\mathbf{Q}^{*}(\mathbf{h}{},\mathbf{a})$ is the value of choosing action $\mathbf{a}$ at history $\mathbf{h}{}$ and then continuing with the optimal policy, $\mathbf{\pi}^{*}$.